# Thunder-NUBench: A Benchmark for LLMs’ Sentence-Level Negation Understanding

### v1

@@ Appendix A Copular Verbs
Table 8: Details of Wikipeda Summary dataset structure with examples.
Appendix B HoVer Dataset
HoVer is constructed on top of the HotpotQA dataset, which is designed to evaluate multi-hop reasoning in question answering. HotpotQA itself is a large-scale collection of Wikipedia-based QA pairs created to address the limitations of prior QA datasets, which often fail to require complex reasoning or explanatory answers (Yang et al., 2018). The construction of HoVer involves rewriting HotpotQA question-answer pairs into claim statements, which are then validated and labeled by annotators. Claims are extended to require multi-hop evidence from up to four Wikipedia articles and are systematically modified to increase complexity. Final labels are assigned as SUPPORTED or NOT-SUPPORTED (Jiang et al., 2020).
Appendix C Wikipedia Summary Dataset
@@
The Wikipedia Summary Dataset contains the titles and introductory summaries of English Wikipedia articles, extracted in September 2017. A summary or introduction in this context refers to the content from the article title up to the content outline (i.e., before the first section heading). The dataset was originally released via GitHub [^3] https://github.com/tscheepers/Wikipedia-Summary-Dataset, but is now accessible through the Hugging Face Hub [^4] https://huggingface.co/datasets/jordiclive/wikipedia-summary-dataset. The dataset license is not explicitly mentioned, but as the original Wikipedia data is distributed under the CC BY-SA 4.0, it is assumed that the dataset would be distributed under the same license. For licensing details, refer to the Wikimedia Terms of Use  [^5] https://foundation.wikimedia.org/wiki/Policy:Terms_of_Use. Table 8 offers an overview of the dataset’s structure. The dataset comprises approximately 430,000 articles, only providing the training set (Scheepers, 2017).
Appendix D Codes for Data Generation
D.1 Sentence-Negation Pair Dataset
@@
To construct the sentence-negation pair dataset, we leverage OpenAI GPT models (OpenAI, 2025) for both grammar correction and sentence merging. We select different model versions depending on the complexity of each task. For sentence merging, which demands nuanced contextual understanding and complex syntax, we use GPT-4. For grammar correction, where edits are more straightforward, GPT-3.5 is sufficient.
@@
Listing 2: Fixing Grammar with OpenAI API.
@@
Listing 3: Merging Sentences with OpenAI API.
D.2 Multiple Choice Dataset
To generate the multiple choice dataset, we first split the large text into single sentences using Python and then automatically generate contradictions and paraphrases for each sentence via the OpenAI API (GPT-4o), followed by human review. The following scripts illustrate the procedures.
@@ D.2 Multiple Choice Dataset
15df = df[df[’sentence’].apply(lambda x: len(x.split()) >= 30)]
@@
Listing 4: Sentence extraction and preprocessing from Wikipedia summaries.
@@ 31 Contradicted sentence:
Listing 5: Contradiction generation with GPT-4o.
@@ 8 Paraphrased sentence:
Listing 6: Paraphrase generation with GPT-4o.
Appendix E Detailed Principles and Examples of the Thunder-NUBench
Standard negation refers to the process of reversing the truth value of the main predicate of a sentence, typically by directly negating the main verb in the main clause, or (in some cases) by using a complementary antonym. The goal is to deny the entire proposition while preserving all other elements of the sentence (subject, objects, temporal context, etc.) unchanged.
Paraphrasing before Negation.
Negation of Simple Sentences.
Negation in Compound Sentences.
When multiple clauses or propositions are coordinated (e.g., with "and", "or", "but"), standard negation is logically applied, governed by De Morgan’s laws:
  • For statements of the form "A and B", the negation is "$\neg$¬A or $\neg$¬B".
  • For statements of the form "A or B", the negation is "$\neg$¬A and $\neg$¬B".
For example, "He passed the test and received an award." $\rightarrow$ "He did not pass the test or did not receive an award."
For compound structures involving three or more elements, the same logic applies:
  • A, B, and C = (A, and B), and C
    $\rightarrow$ $\neg$(A, and B), or $\neg$¬C
    = ($\neg$¬A, or $\neg$¬B), or $\neg$¬C
    = $\neg$¬A, $\neg$¬B, or $\neg$¬C
  • A, B, or C = (A, or B), or C
    $\rightarrow$ $\neg$(A, or B), and $\neg$¬C
    = ($\neg$¬A, and $\neg$¬B), and $\neg$¬C
    = $\neg$¬A, $\neg$¬B, and $\neg$¬C
  • A and B or C = (A, and B), or C
    $\rightarrow$ $\neg$(A, and B), and $\neg$¬C
    = ($\neg$¬A, or $\neg$¬B), and $\neg$¬C
    = $\neg$¬A, or $\neg$¬B, and $\neg$¬C
When strict application of logical negation produces unnatural language, sentences may be split or slightly rephrased for fluency, provided logical meaning is preserved. For example,
  • Original: "He finished the report and submitted the assignment."
Coordinated Elements in the Sentence.
  • If the coordination introduces semantically distinct propositions, that is, each conjunct could form a complete sentence on its own, negation must be applied to each proposition individually. For example,
    "My sister and I studied hard."
  • Conversely, if the coordination connects elements that jointly participate in a single action or state (e.g., a shared subject or a collective predicate), then the sentence is treated as a simple clause, and the predicate as a whole is negated. Logical decomposition is not appropriate. For example,
    "My sister and I share clothes."
    (NOT: "My sister does not share clothes, or I do not share clothes.")
  • This distinction is crucial: even if two noun phrases are coordinated, if the sentence semantically decomposes into separate propositional content, standard negation must apply to each sub-proposition. Otherwise, it applies to the whole predicate as one unit.
  • Other examples of semantically collective predicates where logical splitting is not appropriate include: "be the same", "have in common", "do something together", "combine", "unite", etc. These describe inherently joint or relational properties, not independent propositions. For example,
    "Clarence Brown and Peter Glenville are from the same country." $\rightarrow$ "Clarence Brown and Peter Glenville are not from the same country."
Use of Antonyms.
    • alive / dead
    • true / false
    • present / absent
    • occupied / vacant
    • hot / cold
    • happy / sad
    • tall / short
    • young / old
    • parent / child
    • teacher / student
    • buy / sell
    • employer / employee
General Principles of Standard Negation.
  • If the negated clause creates a contradiction with other parts of the sentence, the contradictory clause must be removed. For example,
    "While the spatial size of the entire universe is unknown, it is possible to measure the size of the observable universe, which is approximately 93 billion light-years in diameter."
    $\rightarrow$ "While the spatial size of the entire universe is unknown, it isn’t possible to measure the size of the observable universe."
    (The relative clause must be removed due to semantic contradiction inside one sentence.)
Common Negation Errors and Corrections.
    • Explanation: Two clauses connected by "while" are not coordinated propositions (they are not equally connected), but instead present contrastive information. Therefore, it is incorrect to apply logical negation across both clauses. Only the relevant clause (e.g., the first statement) should be negated.
Appendix F Compound Sentences and Coordinating Conjunction
A compound sentence consists of two or more independent clauses joined by a coordinating conjunction. Each clause can stand alone, but they are combined to express related ideas (Gleitman, 1965).
  • "She studied hard, and she passed the exam."
  • "I wanted to go, but it was raining."
  • "You can call me, or you can send an email."
Only clearly coordinating conjunctions (and, or, but) are considered for clause-level negation in this benchmark, indicating that two or more independent clauses are equally connected.
Appendix G Negation Phenomena Considered but Omitted
G.1 Double Negation
Double negation refers to the use of two forms of grammatical negation within a single sentence. In standard English, only one negative form should be present in a subject-predicate construction; the presence of two negatives is generally considered non-standard and often results in an unintended meaning. For example, while "He’s going nowhere" is correct, "He’s not going nowhere" is ungrammatical. Another example is "I won’t bake no cake," which combines verb negation ("won’t") with object negation ("no cake"), resulting in a grammatically incorrect construction (Déprez et al., 2015).
In English, certain double negation constructions convey affirmative meanings rather than intensifying negation, effectively paraphrasing the original positive statement (e.g., $\neg\neg p\approx p$(Van der Wouden, 1996). This rhetorical device, known as litotes, often manifests in expressions such as "not bad," implying "good," or "not unhappy," implying "happy." Leveraging this phenomenon, we generated paraphrase candidates for our dataset using such double negation patterns. For example,
However, upon closer examination, these paraphrase candidates do not always preserve the exact meaning of the original sentence. The antonyms used (e.g., "exclude" for "cover," "unhappy" for "honoured") are not always true complementary antonyms, which does not effectively negate the meaning. Moreover, the litotes construction ("does not fail to fuse") tends to add an emphatic nuance, rather than being a perfect semantic equivalent. Therefore, the boundary between paraphrasing and double negation is ambiguous, and their relationship requires more careful analysis. Given these issues, and because our primary focus is on standard, direct forms of negation, we ultimately decide to exclude double negation constructions as paraphrase candidates from our dataset.
G.2 Conditional Statement
Negating conditional statements presents additional challenges, as the intuitive approach to negation often diverges from the logical form. For example, the negation of "If I study hard, I will pass the bar exam" is frequently (but incorrectly) interpreted as "If I do not study hard, I will not pass the bar exam."
Letting "I study hard" be A and "I will pass the bar exam" be B, the original conditional "If I study hard, I will pass the bar exam" (A $\Rightarrow$ B) is logically equivalent to $\neg A\lor B$. The incorrect logical negation, "If I do not study hard, I will not pass the bar exam" ($\neg$¬A $\Rightarrow$ $\neg$¬B), corresponds to $A\lor\neg B$. The correct negation should be $\neg A\lor B$rightarrow $A\land\neg B$, which means "I study hard and I will not pass the bar exam" (Nguyen et al., 2023).
Such distinctions reveal that logical negation and natural language intuition often do not align. Furthermore, forming proper paraphrases for the negation of conditionals is not straightforward and frequently requires extensive rewriting rather than a simple insertion of "not" or converting the verb to its complementary antonym. Since our task is centered on the correct application of verbal negation in main predicates, rather than the broader domain of logical negation in complex constructions, we choose to exclude conditional statements from our dataset.
G.3 Local Negation Constructions Excluded
In constructing the Thunder-NUBench dataset, we consider various types of local (i.e., subclausal) negation, where negation applies to a phrase or constituent rather than the main predicate. However, several constructions are excluded due to their semantic ambiguity, syntactic irregularity, or misalignment with the benchmark’s focus on predicate-level negation.
Infinitive Phrase Negation.
Infinitive phrases (e.g., "to go") can be negated with "not" (e.g., "not to go" or "to not go"). Although grammatically correct, this construction is relatively rare and sounds awkward depending on the context.
  • Original: George wants to go to the park.
  • Negated (infinitive): George wants not to go / George wants to not go.
Since our benchmark targets the negation of predicates, both when generating standard negation and local negation, we exclude this type of infinitive phrase negation, which typically affects subordinate or embedded phrases.
Appositive Clause Negation.
  • Original: My brother, a talented musician, plays the guitar.
Such changes alter descriptive content rather than reversing the meaning of the predicate, and often fall into the domain of contradiction. Accordingly, they were excluded from the dataset.
Prepositional Phrase Negation.
  • Original: She went to the park with her bird.
  • Negated (preposition): She went to the park without her bird.
In all of the above cases, the negation does not target the verb but rather peripheral elements within the sentence. As the Thunder-NUBench is designed to evaluate verbal negation, these local or phrase-level forms of negation were intentionally left out.
Appendix H Thunder-NUBench Structure
@@
Sentence-Negation Pair Dataset.
@@
Multiple-Choice Dataset.
@@
Local Negation Types and Distributions.
choice2_type definition Validation Set Test Set
@@
30 290
compound_part negation applied to one clause within a compound sentence. 29 277
@@
8 105
@@
The details of choice2 types and distribution on validation and test sets are described in Table 9.
Appendix I Models We Use in Experiments
We evaluate two primary model size groups: 3–4 billion (B) parameters and 7–8 billion (B) parameters. Each group contains both pretrained models and instruction-tuned models, as detailed below (Qwen et al., 2025; Grattafiori et al., 2024; Team et al., 2025; Jiang et al., 2023; Team et al., 2024):
  • 3-4B models:
    • Pretrained: Llama-3.2-3B, Qwen2.5-3B, gemma-3-4b-pt
    • Instruction-tuned: Llama-3.2-3B-Instruct, Qwen2.5-3B-Instruct, gemma-3-4b-it
  • 7-8B models:
    • Pretrained: Mistral-7B-v0.3, Llama-3.1-8B, Qwen2.5-7B, gemma-7b
    • Instruction-tuned: Mistral-7B-Instruct-v0.3, Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct, gemma-7b-it
Appendix J Prompt Selection for In-Context Learning
prompt index prompt Type Detail
prompt 1 Default
Negate the following sentence using standard logical negation:
{doc[’sentence’]}
Negated:
prompt 2 Default
Logically negate the following statement:
{doc[’sentence’]}
Negation:"
prompt 3 Default
Reverse the truth value of the statement.
Sentence: {doc[’sentence’]}
Answer:
prompt 4 Default
Reverse the entire meaning of the following sentence using logical and standard negation.
Sentence: {doc[’sentence’]}
Negated:
prompt 5 Definition-based + Default
Standard negation involves inserting ’not’ or using a complementary antonym. Logically negate:
{doc[’sentence’]}
Negation:
prompt 6 Structured + Definition-based
Standard negation involves negating a verbal declarative sentence either by inserting ’not’ after the
primary or auxiliary verb, or by replacing a predicate with its complementary antonym of the main clause.
Logically apply the correct standard negation for the following sentence: {doc[’sentence’]}
Standard negation:
prompt 7 structured + instruction-based
Logically negate the sentence below. If the sentence includes ’A and B’, use ’not A or not B’. If it includes ’A or B’,
use ’not A and not B’. Also apply ’not’ or use complementary antonyms on the main verb(s) of the entire sentence.
Sentence: {doc[’sentence’]}
Negation:"
prompt 8 structured + instruction-based
Apply standard logical negation to the following sentence. This includes:
1. Syntactic negation using ’not’ or auxiliaries on the main verb.
2. Logical negation (e.g., using De Morgan’s laws: ’A and B’ becomes ’not A or not B’).
3. Lexical negation using complementary antonyms on the main predicate (e.g., ’alive’ → ’dead’).
Sentence: {doc[’sentence’]}
Negated version:"
prompt 9
structured + definition,
(comprehensive) instruction-based
Standard negation refers to the basic structural way of negating declarative verbal main clauses in English.
Apply standard logical negation to reverse the truth value of the following sentence.
This includes negating the main predicate by applying:
1. Syntactic negation using ’not’ or auxiliaries.
2. Logical negation (e.g., using De Morgan’s laws: ’A and B’ becomes ’not A or not B’).
3. Lexical negation using complementary antonyms (e.g., ’alive’ → ’dead’).
Sentence: {doc[’sentence’]}
Negated:
prompt 10 role-based + default You are a logic expert. Provide the logically correct standard negation of the following sentence: ’{doc[’sentence’]}’
prompt 11 Reasoning-based
Think step-by-step to logically negate the following sentence. Use standard negation: syntactic negation (’not’)
or complementary antonym on main predicate, and logical rules (e.g., De Morgan’s law).
Reverse the truth value of the statement.
Sentence: {doc[’sentence’]}
Negated version:"
prompt 12 Reasoning-based
Let’s negate this sentence step by step:
1. Identify the core propositions.
2. Apply logical negation to each part.
3. Construct the final negated sentence.
Sentence: {doc[’sentence’]}
Negation:
Table 10: Overview of the prompt variants used for in-context learning in Thunder-NUBench experiments. Prompt 7 is the base prompt of the whole paper. Prompts 1, 7, and 11 are used in experiments (each as default, instruction-based, and reasoning-based prompt).
We explore a range of prompt types, from basic to structured and reasoning-based instructions (Zhao et al., 2021; Li, 2023; Wan et al., 2023). Table 10 shows the details of all prompt templates explored. Among the prompt variants, prompt 1 (default), prompt 7 (instruction-based), and prompt 11 (reasoning-based) were selected for use in the experiments. Only the best-performing prompt on the validation set is reported in the main text (Section 5).
J.1 Prompt Type Design
Llama gemma Qwen
Llama-3.2-3B Llama-3.2-3B-Instruct gemma-3-4b-pt gemma-3-4b-it Qwen2.5-3B Qwen2.5-3B-Instruct Average
prompt
index
prompt
type
acc acc_norm acc acc_norm acc acc_norm acc acc_norm acc acc_norm acc acc_norm acc acc_norm
prompt 1 0.550 0.550 0.590 0.600 0.370 0.410 0.400 0.430 0.580 0.590 0.590 0.640 0.513 0.537
prompt 2 0.520 0.510 0.600 0.620 0.360 0.420 0.400 0.400 0.540 0.570 0.600 0.620 0.503 0.523
prompt 3 0.430 0.440 0.520 0.520 0.350 0.380 0.380 0.410 0.460 0.490 0.490 0.510 0.438 0.458
prompt 4 0.500 0.550 0.590 0.620 0.440 0.470 0.440 0.460 0.530 0.570 0.590 0.640 0.515 0.552
prompt 5 default 0.480 0.510 0.470 0.490 0.380 0.400 0.420 0.460 0.510 0.520 0.560 0.580 0.470 0.493
prompt 6 0.540 0.560 0.620 0.650 0.440 0.480 0.430 0.460 0.520 0.550 0.590 0.610 0.523 0.552
prompt 7 0.530 0.540 0.590 0.620 0.470 0.510 0.460 0.460 0.590 0.610 0.770 0.770 0.568 0.585
prompt 8 0.510 0.510 0.580 0.600 0.510 0.520 0.500 0.520 0.580 0.600 0.610 0.660 0.548 0.568
prompt 9 instruction 0.530 0.530 0.580 0.610 0.500 0.530 0.540 0.550 0.580 0.570 0.600 0.630 0.555 0.570
prompt 10 role 0.450 0.480 0.410 0.440 0.360 0.420 0.450 0.510 0.550 0.560 0.620 0.660 0.473 0.512
prompt 11 0.570 0.590 0.620 0.640 0.440 0.450 0.530 0.540 0.580 0.590 0.610 0.640 0.558 0.575
prompt 12 reasoning 0.550 0.550 0.630 0.670 0.450 0.460 0.510 0.510 0.600 0.600 0.600 0.630 0.557 0.570
Table 11: Prompt-wise Results on the Validation Set for 3B–4B Models. The red values indicate the highest performance in each column. acc denotes the accuracy, and acc_norm indicates the length-normalized accuracy.
Mistral Llama gemma Qwen
Mistral-7B-v0.3 Mistral-7B-Instruct-v0.3 Llama-3.1-8B Llama-3.1-8B-Instruct gemma-7b gemma-7b-it Qwen2.5-7B Qwen/Qwen2.5-7B-Instruct average
prompt
index
prompt
type
acc acc_norm acc acc_norm acc acc_norm acc acc_norm acc acc_norm acc acc_norm acc acc_norm acc acc_norm acc acc_norm
prompt 1 0.540 0.540 0.620 0.660 0.490 0.520 0.610 0.640 0.510 0.540 0.630 0.690 0.630 0.650 0.680 0.700 0.589 0.618
prompt 2 0.480 0.500 0.570 0.630 0.510 0.520 0.560 0.590 0.520 0.560 0.620 0.630 0.630 0.660 0.670 0.670 0.570 0.595
prompt 3 0.420 0.460 0.520 0.560 0.480 0.520 0.500 0.510 0.500 0.520 0.540 0.570 0.470 0.510 0.540 0.540 0.496 0.524
prompt 4 0.430 0.470 0.610 0.650 0.450 0.460 0.520 0.570 0.500 0.530 0.540 0.570 0.540 0.550 0.650 0.670 0.530 0.559
prompt 5 default 0.490 0.520 0.710 0.710 0.460 0.470 0.450 0.470 0.500 0.540 0.560 0.580 0.530 0.550 0.560 0.600 0.533 0.555
prompt 6 0.520 0.550 0.690 0.730 0.490 0.510 0.560 0.590 0.520 0.530 0.520 0.570 0.540 0.570 0.520 0.570 0.545 0.578
prompt 7 0.540 0.570 0.740 0.750 0.500 0.520 0.670 0.670 0.570 0.570 0.640 0.640 0.600 0.620 0.670 0.710 0.616 0.631
prompt 8 0.530 0.540 0.700 0.730 0.470 0.500 0.570 0.600 0.540 0.540 0.720 0.720 0.640 0.660 0.660 0.690 0.604 0.623
prompt 9 instruction 0.570 0.580 0.730 0.730 0.460 0.490 0.560 0.600 0.540 0.550 0.610 0.620 0.600 0.600 0.650 0.680 0.590 0.606
prompt 10 role 0.500 0.520 0.630 0.650 0.460 0.480 0.470 0.540 0.500 0.520 0.660 0.750 0.610 0.630 0.620 0.640 0.556 0.591
prompt 11 0.520 0.550 0.650 0.660 0.510 0.530 0.550 0.610 0.580 0.610 0.710 0.740 0.670 0.690 0.680 0.690 0.609 0.635
prompt 12 reasoning 0.540 0.540 0.680 0.720 0.530 0.550 0.610 0.610 0.580 0.570 0.630 0.650 0.590 0.620 0.590 0.620 0.594 0.610
Table 12: Prompt-wise Results on the Validation Set for 7B–8B Models. The red values indicate the highest performance in each column. acc denotes the accuracy, and acc_norm indicates the length-normalized accuracy.
Analysis of prompt type validation results reveals that prompt 7, which provides an instruction-based, structured formulation, achieves the highest overall performance. Based on these results, we adopt this prompt as the base template for our main experiments. Additionally, among the default-type prompts, prompt 1 shows the best average performance for 7–8B models, while among reasoning-based prompts, prompt 11 shows better performance among all models. See the details in Table 11 and Table 12. Therefore, we use these prompts as the default prompt (prompt 1), instruction prompt (prompt 7), and reasoning prompt (prompt 11) in the experiment results of Appendix M.
The optimal prompt type or format can differ depending on model series and size, highlighting that the choice of prompt significantly impacts model performance. Nevertheless, detailed instructions or reasoning-based formats generally outperform minimal "default" prompts, underscoring the importance of providing explicit context and guidance in in-context learning settings.
J.2 Prompt Wording Variants
3-4B models average 7-8B models average
prompt index Detail acc acc_norm acc acc_norm
[…]
Table 13: Prompt Wording Variants Used for Validation Set Evaluation. The base prompt is prompt 7 from Table 10. The red values indicate the highest performance in each column. acc denotes the accuracy, and acc_norm indicates the length-normalized accuracy.
We further investigate the effect of prompt wording by testing minor variants in the phrasing of instructions and output cues, while maintaining the same underlying task specification. For example, all prompts in this set share the same core instruction,
"Logically negate the sentence below. If the sentence includes ’A and B’, use ’not A or not B’. If it includes ’A or B’, use ’not A and not B’. Also apply ’not’ or use complementary antonyms on the main verb(s) of the entire sentence."
of the prompt 7, but differ in the wording of the instruction or the label for the output (e.g., "Negation:", "Negated:", "Output:").
See Table 13 for a full list of variants. Even such superficial changes in wording can affect performance, further emphasizing the importance of careful prompt engineering.
[…]
Appendix K Evaluating via Language Model Evaluation Harness
This section describes how Thunder-NUBench and ScoNe is integrated into the LM Evaluation Harness (Gao et al., 2024) for zero-shot, few-shot evaluation. We use the prompt for ScoNe according to the original paper (Hypothesis Question Prompt format) (She et al., 2023).
1task: nubench
@@ Appendix I Models We Use in Experiments
3dataset_name: null
@@
5validation_split: validation
6test_split: test
7fewshot_split: validation
8process_docs: !function utils.process_docs
9doc_to_text: "{{query}}"
10doc_to_target: "{{gold}}"
11doc_to_choice: "choices"
12metric_list:
13 - metric: acc
14 aggregation: mean
15 higher_is_better: true
16 - metric: acc_norm
17 aggregation: mean
18 higher_is_better: true
Listing 7: NUBench/NUBench.yaml
1import re
2import datasets
3
4def process_docs(dataset: datasets.Dataset) -> datasets.Dataset:
5 def _process_doc(doc):
[…]
7
8 if doc.get("choice2_type", "") == "non-applicable":
9 choices = [doc["choice1"], doc["choice3"], doc["choice4"]]
10 else:
12
13 return {
14 "query": prompt,
15 "choices": choices,
16 "gold": 0 # always choice1
17 }
18 return dataset.map(_process_doc)
Listing 8: NUBench/utils.py
1task: ScoNe
2dataset_path: tasksource/ScoNe
3dataset_name: null
4output_type: multiple_choice
5training_split: train
6test_split: test
7process_docs: !function utils.process_docs

### v3

@@ Appendix G HoVer Dataset
HoVer is constructed on top of the HotpotQA dataset, which is designed to evaluate multi-hop reasoning in question answering. HotpotQA itself is a large-scale collection of Wikipedia-based QA pairs created to address the limitations of prior QA datasets, which often fail to require complex reasoning or explanatory answers (yang2018hotpotqa). The construction of HoVer involves rewriting HotpotQA question-answer pairs into claim statements, which are then validated and labeled by annotators. Claims are extended to require multi-hop evidence from up to four Wikipedia articles and are systematically modified to increase complexity. Final labels are assigned as SUPPORTED or NOT-SUPPORTED (jiang2020hover).
Appendix H Wikipedia Summary Dataset
@@
Table 9: Details of Wikipedia Summary dataset structure with examples.
@@
The Wikipedia Summary Dataset contains the titles and introductory summaries of English Wikipedia articles, extracted in September 2017. A summary or introduction in this context refers to the content from the article title up to the content outline (i.e., before the first section heading). The dataset was originally released via GitHub [^2] https://github.com/tscheepers/Wikipedia-Summary-Dataset, but is now accessible through the Hugging Face Hub [^3] https://huggingface.co/datasets/jordiclive/wikipedia-summary-dataset. The dataset license is not explicitly mentioned, but as the original Wikipedia data is distributed under the CC BY-SA 4.0, it is assumed that the dataset would be distributed under the same license. For licensing details, refer to the Wikimedia Terms of Use  [^4] https://foundation.wikimedia.org/wiki/Policy:Terms_of_Use. Table 9 offers an overview of the dataset’s structure. The dataset comprises approximately 430,000 articles, only providing the training set (scheepers2017compositionality).
Appendix I Human Review Protocol
To ensure high-quality data construction, we implement a rigorous quality control protocol that combines generation, independent review, and iterative consensus building. The process involves the following key steps:
  • Task allocation and independence. Authors are assigned distinct portions of the dataset, but no author is permitted to review the data they have generated. This ensures that each instance is subject to at least one independent review.
  • Sequential authoring across choices. For the multiple-choice dataset, construction proceeds in four stages: standard negation, local negation, contradiction, and paraphrase. At each stage, different authors are responsible for creating the new option, while reviewers who have not authored that option perform the verification.
  • Cross-checking and layered review. Each newly created option is reviewed by at least one other author, and reviewers also revisit earlier options in the same instance. For example, when reviewing the paraphrased sentence, the reviewer also checks that standard negation, local negation, and contradiction sentences are correct. As a result, every instance undergoes multiple rounds of verification across stages, such that all authors ultimately examine data they have not created themselves.
  • Guideline refinement and retroactive correction. Generation and reviewing guidelines are continuously updated based on discussion of ambiguous or problematic cases. Whenever the guidelines changes, all previously created data are revisited to ensure compliance, promoting consistency across the dataset.
  • Consensus and adjudication. Disagreements are discussed in weekly meetings and, if necessary, adjudicated by a lead reviewer, ensuring that no instance remains unresolved.
Overall, this iterative and layered procedure ensures that every instance in the multiple-choice dataset is independently reviewed multiple times, leading to stable guidelines and a consistent dataset.
Appendix J Detailed Principles and Examples of the Thunder-NUBench
While the main text already defines the core notions of standard and local negation (Section 3) and explains how they are applied throughout dataset construction (Section 4), here we provide more detailed illustrations.
Paraphrasing before Negation.
Negation of Simple Sentences.
Negation in Compound Sentences.
When multiple clauses or propositions are coordinated (e.g., with "and", "or", "but"), standard negation is logically applied, governed by De Morgan’s laws. Here, "but" is treated as a coordinating conjunction equivalent to "and" in terms of logical structure, so its negation follows the same principle.
  • Conjunction "$P$ and/but QQ": the negation is $\text{Neg}(P)\;\text{or}\;\text{Neg}(Q)$".
  • Disjunction "$P$ or QQ": the negation is $\text{Neg}(P)\;\text{and}\;\text{Neg}(Q)$".
For example, "He passed the test and received an award." is negated as "He did not pass the test or did not receive an award."
When application of logical negation produces unnatural language, sentences may be split or slightly rephrased for fluency, provided logical meaning is preserved. For example,
  • Original: "He finished the report and submitted the assignment."
Coordinated Elements in the Sentence.
  • If the coordination introduces semantically distinct propositions, that is, each conjunct could form a complete sentence on its own, negation must be applied to each proposition individually. For example, "My sister and I studied hard."
  • Conversely, if the coordination connects elements that jointly participate in a single action or state (e.g., a shared subject or a collective predicate), then the sentence is treated as a simple clause, and the predicate as a whole is negated. Logical decomposition is not appropriate. For example, "My sister and I share clothes."
    (NOT: "My sister does not share clothes, or I do not share clothes.")
  • This distinction is crucial: even if two noun phrases are coordinated, if the sentence semantically decomposes into separate atomic propositions, standard negation must apply to each atomic proposition. Otherwise, it applies to the whole predicate as one unit.
  • Other examples of semantically collective predicates where logical splitting is not appropriate include: "be the same", "have in common", "do something together", "combine", "unite", etc. These describe inherently joint or relational properties, not independent propositions. For example, "Clarence Brown and Peter Glenville are from the same country." should be negated as "Clarence Brown and Peter Glenville are not from the same country."
Use of Antonyms.
    • alive / dead
    • true / false
    • present / absent
    • occupied / vacant
    • hot / cold
    • happy / sad
    • tall / short
    • young / old
    • parent / child
    • teacher / student
    • buy / sell
    • employer / employee
General Principles of Standard Negation.
  • If the negated clause creates a contradiction with other parts of the sentence, the contradictory clause must be removed. For example, the standard negation of the sentence "While the spatial size of the entire universe is unknown, it is possible to measure the size of the observable universe, which is approximately 93 billion light-years in diameter."
    will be "While the spatial size of the entire universe is unknown, it isn’t possible to measure the size of the observable universe."
    The relative clause must be removed because its content directly contradicts the negated main clause.
Common Negation Errors and Corrections.
    • Explanation: Two clauses connected by while are not coordinated propositions (as with and or or), but instead express contrastive information. Therefore, applying logical negation across both clauses is incorrect. Negation should apply only to the main clause (here, the first statement), while the contrasting clause remains affirmative.
Appendix K Code for Data Construction
K.1 Sentence-Negation Pair Dataset
To construct the sentence-negation pair dataset, we begin by randomly sampling sentences labeled as "supported facts" from the HoVer dataset. Since the original data often contains grammatical errors, we utilize OpenAI’s API (OpenAI) to automatically correct these issues. In cases where the selected text consists of multiple sentences, we merge or split them as needed to ensure that each example is a single sentence, aligning with our sentence-level task objective.
@@ Appendix I Human Review Protocol
We select different model versions depending on the complexity of each task. For sentence merging, which demands nuanced contextual understanding and complex syntax, we use GPT-4. For grammar correction, where edits are more straightforward, GPT-3.5 is sufficient.
@@
Listing 1: Fixing Grammar with OpenAI API.
@@
Listing 2: Merging Sentences with OpenAI API.
K.2 Multiple Choice Dataset
To construct the multiple-choice dataset, we first segment the "summary" column of the Wikipedia Summary dataset, which often contains multiple sentences in a single entry, into individual sentences. To focus on the challenges of negation in complex sentences, we filter out sentences that are too short. This process is done with Python code.
Since conditional sentences (e.g., "If $P$, QQ") are rarely present in the Wikipedia summary dataset, we adopt a two-step approach: (1) prompting the model to generate conditional variants from given sentences (using OpenAI API, GPT-4o-mini), and (2) manually filtering or lightly editing the results to obtain valid conditionals.
Subsequently, we automatically generate contradictions and paraphrases for each sentence via the OpenAI API (GPT-4o) as well, followed by human review. The following scripts illustrate the procedures.
@@
15df = $df[$’sentence’].apply(lambda x: len(x.split()) >= 30)]
@@
Listing 3: Sentence extraction and preprocessing from Wikipedia summaries.
1def generate_conditionals(sentence):
2 prompt = f"""
3 Based on the sentence below, write a conditional sentence that uses the main topic of the sentence.
4 The conditional sentence should express a hypothetical situation or cause-effect relationship related to the topic. It can be slightly complex in structure.
5 For example:
6 - If it rains tomorrow, I will stay home.
7 Sentence:
8 ’{sentence}’
9 """
10
11 completion = client.chat.completions.create(
12 model="gpt-4o-mini",
13 messages=[
14 {"role": "system", "content": "You are a helpful assistant that specializes in generating conditional sentences."},
15 {"role": "user", "content": prompt}
16 ]
17 )
18
19 return completion.choices[0].message.content
Listing 4: Conditionals sentence generation.
@@ 31 Contradicted sentence:
Listing 5: Contradiction generation.
@@ 8 Paraphrased sentence:
Listing 6: Paraphrase generation.
Appendix L Thunder-NUBench Dataset Structure
@@
L.1 Sentence-Negation Pair Dataset
@@
L.2 Multiple-Choice Dataset
@@
choice2_type Definition
Demonstration
Set
Test Set
@@
12 308
@@
12 310
compound_part negation applied to one clause within a compound sentence. 12 294
@@
The details of choice2_type and distribution on demonstration and test sets are described in Table 10. It follows the definition in Table 3.
In addition to the Wikipedia Summary dataset, we supplement the evaluation set with conditionals (e.g., If $P$, $Q$) by manually searching Wikipedia articles where such constructions are more likely to occur (e.g., Newton’s laws of motion). Among the 100 conditional sentences included across demonstration and test sets, 20 are collected through manual search (marked with indices beginning with "S" in wikipedia_index). Meanwhile, the remaining 80 are sampled from the Wikipedia Summary dataset and converted into conditional form using the script in 4 (Appendix K).
Appendix M Incorporating Thunder-NUBench into LM Evaluation Harness
This section describes how Thunder-NUBench is integrated into the LM Evaluation Harness (eval-harness) for zero-shot and few-shot evaluation, both in completion-based and option-selection settings.
1task: nubench_completion
@@
3dataset_name: multiple-choice
@@
5test_split: test
6fewshot_split: demonstration
7process_docs: !function utils.process_docs_completion